gough.ath.cx Homepage v2

This is the personal server of Gough, served from Home.
This is what I mean by a redesign ...

New and Featured Content Radio Related Computer Related Totally Unrelated!
  • This website is now considered a legacy site. No further updates will be committed.

Photos Sefton High Stuff APC Stuff Archives

Site Update Log

13th Feb 2011: Finished week 11 log, draft final report uploaded. Updated excel spreadsheet due to an error which was spotted.
10th Feb 2011: Final poster uploaded. Partial Week 11 ToR log online. Excel spreadsheet of data summary online as well.
5th Feb 2011: Updated Week 10 log for ToR, added v2 of the Poster.
31st Jan 2011: Finalized Week 9 log for ToR, added plot all for oudoor means and all modes.
23rd Jan 2011: Updated Week 8 log for ToR, added partial Week 9.
19th Jan 2011: Updated Week 8 log for ToR.
16th Jan 2011: Updated Taste of Research section with logbook for Week 7 and the beginning of Week 8. Plotallindoor.m and fig now completed. Nice and messy!
9th Jan 2011: Updated Taste of Research section with new logbook entries for Week 6. More pretty graphs and a little analysis - progress is going fairly well. Plus, see my face swell! Updated mX Archives with 5 more October 2010 papers. A bit late, I know, but I've been very busy and didn't have much time to maintain the website.
5th Dec 2010: Taste of Research Engineer's Log Added. You can read about my work with the University at the moment over the holidays in regards to Wi-Fi if you are so inclined.
4th Dec 2010: System was restored to normal after being migrated due to hard disk failure. The previous Maxtor Diamondmax8 Plus 6E040L0 has now been replaced with an IBM Deathstar DPTA-372050 20.5Gb dinosaur, system reinstalled and bought back online. Inbetween, another server took over the role of webserving, but the liev streams were taken offline due to this problem. Uptime is now sub 100 percent - but not due to this incident. Earlier (about five days prior), we suffered a neighbourhood wide power outage which knocked everything out.
31st Oct 2010: Added GPS Chipset Shootout and BD-R's - can you trust them?
29th Oct 2010: Added Chinese Batteries Exposed. Updated the QRSS pages with the experiments being concluded for the time being.
26th Oct 2010: Added QRSS Monitoring to the website. Haven't had much time to do many changes unfortunately.
16th Sept 2010: Added two old papers to the mX Archives.
14th Sept 2010: Site Update Log Added to site to help track updates from now onward. Added two new papers to the mX Archives.

Feel free to contact me at gough.beta(removethis)@yahoo.com.au
With thanks to:

TPG Internet.